OpenStudy (anonymous):

A sector has an area of 942.5cm^2 and an arc length of 62.8cm what is the radius of the circle??

sam (.sam.):

Arc length \[s=r \theta\] \[62.8=r \theta\] \[r=\frac{62.8}{\theta}\] ------------------------- \[Area =\frac{1}{2}r^2 \theta\] \[\frac{1}{2}(\frac{62.8}{\theta})^2 \theta=942.5\] \[\theta=2.09~ radians\] r=30

OpenStudy (anonymous):

did you use a calculator of the computations?

sam (.sam.):

For the formula, no For the calculations for the answer, yes

OpenStudy (anonymous):

how would I isolate the θ??

sam (.sam.):

\[\frac{1}{2}(\frac{62.8}{\theta})^2 \theta=942.5\] multiply 2 both sides \[(\frac{62.8}{\theta})^2 \theta=1885\] \[(\frac{62.8^2}{\theta^2}) \theta=1885\] \[(\frac{62.8^2}{\theta^\cancel2}) \cancel\theta=1885\] \[(\frac{62.8^2}{\theta}) =1885\] \[(\frac{62.8^2}{1885}) =\theta\]

OpenStudy (anonymous):

Thank you so very much.